# SoTA with Less: MCTS-Guided Sample Selection for Data-Efficient Visual Reasoning Self-Improvement

### v1

@@ 3.3 Visual Reasoning Training
Table 3: Visual reasoning training data comparison between ThinkLite-VL and other VLM reasoning models. ALL these reasoning models have distilled knowledge from larger models or closed-source models except for MM-Eureka-Qwen-7B. MM-Eureka-Qwen-7B uses more K12 data (54k) than ours and performs accuracy-based data filtering before training. Here the data size refers to the amount of additional visual reasoning data used to boost the base model for reasoning via SFT or RL training.
@@
After conducting MCTS-based sample selection and obtaining a filtered set of 11k high-quality training data, we then perform RL fine-tuning on the Qwen2.5-VL-7B-Instruct model using these selected data. Specifically, we employ Group Relative Policy Optimization (GRPO) loss function proposed by [60] for training, with the objective defined as follows:
@@
We provide the training prompt template during RFT in Table 4.
[…]
@@ 4.1 Benchmark Evaluation
In this subsection, we systematically evaluate ThinkLite-VL on several commonly used multimodal benchmark datasets and perform comprehensive comparisons with existing reasoning models. Through these experiments, we demonstrate the effectiveness and advantages of our model in multimodal reasoning tasks.
Baselines and implementation details
We use Qwen2.5-VL-7B-Instruct as the base model and perform RFT on the 11k high-quality data obtained through MCTS-based filtration, resulting in our proposed model, named ThinkLite-VL-7B. We conduct training using Easy-R1 [97] code base and set GRPO rollout number as 32. Our main baselines are as follows: (1) Qwen2.5-VL-7B-Instruct, serving as our base model; (2) ThinkLite-VL-Random11k, trained using RFT on a randomly sampled subset of 11k instances from the full dataset. Besides, we report the performance of several recent reasoning VLMs for comparison, including the SFT-based models LLaVA-Cot-11B and Mulberry-7B, as well as the RFT-based models Vision-R1, MM-Eureka-Qwen-7B, and OpenVLThinker-7B. We also include larger open-source models and commercial models as SOTA performance references which include Qwen2.5-VL-72B-Instruct, InternVL2.5-78B, GPT-4o, and O1.
@@
Table 5: Comparison of different VLMs on 8 widely used visual benchmarks. The grey sections indicate models with larger parameter sizes and closed-source models. Our model achieves SoTA performance at the 7B level on 6 benchmarks and reaches a SoTA performance of 75.1 on MathVista among all VLMs. On average, our model improves performance by 7% compared with Qwen2.5-VL-7B-Instruct. We do not evaluate Mulberry-7B on MathVision because Mulberry-7B uses MathVision as training dataset, and for Vision-R1-7B, their model is not open-sourced, so we only refer to the results reported in their paper.
@@
Qwen2.5-VL-72B-Instruct 74.8 39.8 57.6 70.2 70.8 88.6 76.2 88.5 70.81
InterVL2.5-78B 72.3 34.9 51.7 70.1 69.5 88.3 72.3 89.1 68.53
GPT-4o 63.8 36.8 50.2 69.1 64.7 83.4 69.1 84.6 65.21
O1 73.9 78.2
@@
MM-EUREKA-Qwen-7B 54k 73.0 31.9 50.3 52.3 64.1 79.3 64.9 81.4 62.15
Qwen2.5-VL-7B-Instruct 67.8 23.6 44.5 50.6 61.7 80.7 66.0 82.6 59.69
ThinkLite-VL-7B 11k 75.1 32.9 50.7 54.6 65.0 81.4 67.8 83.6 63.89
SoTA performance over 7B reasoning models
@@
As shown in Table 5, ThinkLite-VL-7B shows a significant improvement in average performance across the eight benchmarks compared to the base model Qwen2.5-VL-7B-Instruct, with the average performance increasing from 59.69 to 63.89. Compared to ThinkLite-VL-Random11k, which is trained with the same data size using random sampling, our method shows significant advantages across all benchmarks, indicating the effectiveness and importance of MCTS-based sample selection. Furthermore, ThinkLite-VL-7B also outperforms reasoning models that primarily achieve performance enhancement through extensive knowledge distillation (such as LLaVA-CoT-11B, Mulberry-7B, Vision-R1-7B, and OpenVLThinker-7B) with the closest average performance to GPT-4o. Compared to MM-EUREKA-Qwen-7B, which does not involve SFT knowledge distillation but adopts a larger RL training dataset, our model consistently outperforms across all benchmarks, highlighting the importance of high-quality data filtering before training, and the effectiveness of the proposed MCTS-based filtering. From the perspective of individual benchmarks, our method achieves the highest scores among 7B-level models on six out of the eight benchmarks. The only exceptions are the MMMU and MathVerse benchmarks, where we slightly lag behind Mulberry-7B and Vision-R1-7B that focused on a narrower range of tasks, respectively. Remarkably, our model achieves the SoTA accuracy of 75.1 on the MathVista benchmark, surpassing larger open-sourced VLMs, GPT-4o, and O1.
4.2 Importance of MCTS-based Sample Selection
In this section, we conduct ablation studies to demonstrate the importance of MCTS-based sample selection. We compare five different training settings of ThinkLite-VL: (1) ThinkLite-VL-Unsolved: Trained using only the 5.6k samples that could not be solved by MCTS, representing the most difficult subset. (2) ThinkLite-VL-Iter5Only: Trained on the subset of data that VLM is able to solve via MCTS, but required more than 5 iterations. This set, combined with the unsolved samples, forms the full 11k training set used in ThinkLite-VL. (3) ThinkLite-VL-Random11k: Trained on a randomly sampled 11k subset from the full 70k dataset, matching the size of the ThinkLite-VL training set. (4) ThinkLite-VL-SelfConsistency: Trained on 23k samples selected based on a self-consistency difficulty measure. Specifically, for each prompt, we perform 50 rollouts using Qwen2.5-VL-7B-Instruct and compute answer accuracy using Qwen2.5-7B-Instruct. Samples with accuracy lower than 0.2 are selected for RFT. (5) ThinkLite-VL-Fullset: Trained on the complete 70k dataset without any filtering. We report the evaluation results of all five settings across the eight VLM benchmarks, as shown in Table 6.
We observe that ThinkLite-VL-7B, trained using 11k samples via MCTS-guided sample selection, achieves the highest average performance (63.89) among all settings. It outperforms not only the random sampling baseline (ThinkLite-VL-Random11k, 60.89) but also models trained on the full dataset (ThinkLite-VL-Fullset, 63.13) and self-consistency-based filtering (ThinkLite-VL-SelfConsistency, 63.15), despite using significantly fewer training samples. This highlights the effectiveness of our difficulty-aware data selection strategy. Further analysis reveals that models trained on subsets derived solely from unsolved samples (ThinkLite-VL-Unsolved, 62.04) or samples requiring more than five iterations (ThinkLite-VL-Iter5Only, 62.38) also show decent performance, suggesting that hard and medium-difficulty samples contribute meaningfully to reasoning ability. However, neither subset alone is sufficient. The combination of both unsolved and medium-difficulty samples yields the strongest and most effective training signal.
Besides, we compare the reward curves during RFT of ThinkLite-VL-Random11k, ThinkLite-VL-Fullset, ThinkLite-VL-Iter5Only, and ThinkLite-VL, as shown in Figure 5. Although ThinkLite-VL-Random11k and ThinkLite-VL-Fullset achieve higher rewards during training, their actual benchmark performances are inferior to ThinkLite-VL. This observation suggests that incorporating a large number of easy samples into training rapidly improves rewards but fails to enhance the model’s reasoning ability. Moreover, ThinkLite-VL exhibits notably lower rewards compared to ThinkLite-VL-Iter5Only, indicating that the unsolved data identified by our MCTS-based sample selection strategy indeed pose significant challenges to the VLM. By progressively learning to solve these challenging problems during training—even if not all are solved completely—the reasoning capabilities of VLMs can be substantially improved.
@@ 4.2 Importance of MCTS-based Sample Selection
Figure 5: Comparison of reward curves of models trained with different data during RFT. Iter5+Unsolved 11k dataset presents the most challenging learning setting for VLM, highlighting the difficulty of the samples selected by MCTS-based sample selection.
4.3 Ablation Study of Data Difficulty
@@
Table 7: ThinkLite-VL performance under different training data difficulty settings. Iter5+Unsolved achieves the best performance.
@@
5 Case Studies
In this section, we present samples of varying difficulty levels selected by the MCTS-based sample selection method across different datasets, as shown in Tables 13 through 12. The difficulty levels are determined based on the number of reasoning iterations required by the VLM to arrive at the correct answer during the MCTS process, providing reference examples for understanding how the method distinguishes between easy and challenging samples.

### v2

@@ 3.3 Visual Reasoning Training
Table 1: Visual reasoning training data comparison between ThinkLite-VL-7B and other 7B-level VLM reasoning models. ALL these reasoning models have distilled knowledge from larger models or closed-source models except for MM-Eureka-Qwen-7B. MM-Eureka-Qwen-7B performs accuracy-based data filtering before training and uses more data (15k) than ours. Here the data size refers to the amount of extra visual reasoning data used to boost the base model for reasoning, via SFT or RFT.
@@
After conducting MCTS-based sample selection and obtaining a filtered set of high-quality training data (11k for 7B and 7.5k for 72B), we then perform RL fine-tuning on the Qwen2.5-VL models using these selected data. Specifically, we employ Group Relative Policy Optimization (GRPO) loss function proposed by [62] for training, with the objective defined as follows:
@@
We provide the training prompt template during RFT in Appendix A Table 8.
@@ 4.1 Benchmark Evaluation
We systematically evaluate ThinkLite-VL on several commonly used multimodal benchmark datasets and perform comprehensive comparisons with existing reasoning models. Through these experiments, we demonstrate the effectiveness and advantages of our model in multimodal reasoning tasks.
Baseline VLMs.
We compare our method with both 7B level and 72B level models as follows:
• For 7b-level VLMs, we use Qwen2.5-VL-7B-Instruct as the base model and perform RFT on the 11k high-quality data obtained through MCTS-based filtration, resulting in our reasoning model, named ThinkLite-VL-7B. We conduct training using Easy-R1 [101] code base and set GRPO rollout number as 32. Our main baselines are as follows: (1) Qwen2.5-VL-7B-Instruct [3], serving as our base model; (2) ThinkLite-VL-Random11k, trained using RFT on a randomly sampled subset of 11k instances from the full 70k dataset. Besides, we report the performance of several recent general and reasoning VLMs for comparison, including general opensourced models LLaVA-Onevision-7B [30] and InternVL2.5-8B [10], the SFT-based reasoning models LLaVA-Cot-11B [83] and Mulberry-7B [86], as well as the RFT-based reasoning models Vision-R1 [22], MM-Eureka-Qwen-7B [50], and OpenVLThinker-7B [12].
• For 72B-level VLMs, we use Qwen2.5-VL-72B-Instruct as the base model. We perform RFT on the 7.5k high-quality data obtained by Qwen2.5-VL-72B-Instruct through MCTS-based filtration and get 72B reasoning model ThinkLite-VL-72B. The 72B-level baselines include: (1) our base model Qwen2.5-VL-72B-Instruct [3]; (2) two opensourced general VLMs LLaVA-Onevision-72B [30] and InternVL2.5-78B [10]; (3) one opensouced reasoning model QvQ-72B [71]; (4) ThinkLite-VL-Random7.5k, trained using RFT on 7.5k randomly selected samples from the full 70k dataset. We also include proprietary models as performance references which include OpenAI-GPT-4o and OpenAI-o1. For all models, we use 8$\times$80G A100 GPUs for model training and evaluation.
@@
Table 2: Comparison of different VLMs on 8 widely used visual benchmarks. Our model achieves SoTA performance at both 7B level and 72B level on 6 benchmarks and reaches a SoTA performance of 79.7 on MathVista among all VLMs. On average, our model improves performance by 7.5% and 6.5% compared with our base models Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct. We do not evaluate Mulberry-7B on MathVision because Mulberry-7B uses MathVision as training dataset. We evaluate all models with same code using vLLM [27] inference. For reasoning models, we use thinking templates provided in their codebase to generate thoughts and get the final answer.
@@
Proprietary Models
OpenAI-GPT-4o 63.8 36.8 50.2 69.1 64.7 83.4 69.1 84.6 65.21
OpenAI-o1 73.9 58.2 57.0 77.6
7B-level General and Reasoning Vision-Language Models
LLaVA-Onevision-7B 63.2 17.4 26.2 48.8 61.7 80.8 57.5 81.4 54.63
InternVL2.5-8B 64.4 22.0 39.5 54.9 62.8 82.7 68.8 83.3 59.80
Qwen2.5-VL-7B-Instruct 67.8 23.6 44.5 50.6 61.7 80.7 66.0 82.6 59.69
@@
MM-EUREKA-Qwen-7B 15k 73.0 31.9 50.3 52.3 64.1 79.3 64.9 81.4 62.15
Our 7B-level Reasoning Model
ThinkLite-VL-7B-Random11k 11k 71.9 26.1 47.3 51.7 62.7 81.1 65.5 80.9 60.89
ThinkLite-VL-7B 11k 75.1 32.9 52.1 55.5 65.0 81.4 67.8 83.6 64.18
$\Delta$ (Ours - Random selection) +3.2 +6.8 +4.8 +3.8 +2.3 +0.3 +2.3 +2.7 +3.29
$\Delta$ (Ours - Open 7B SoTA) +1.6 +1.0 +0.2 +0.5 +0.9 -1.3 -1.0 +0.3 +2.03
72B-level General and Reasoning Vision-Language Models
LLaVA-Onevision-72B 67.5 29.3 39.1 56.8 66.1 85.9 63.7 85.6 61.75
InterVL2.5-78B 72.3 34.9 51.7 68.7 68.9 87.2 72.3 87.9 67.99
Qwen2.5-VL-72B-Instruct 74.8 35.2 53.3 63.4 68.4 87.4 76.3 87.2 68.25
QvQ-72B 71.4 32.7 48.6 70.3 67.2 86.3 75.9 86.6 67.37
Our 72B-level Reasoning Model
ThinkLite-VL-72B-Random7.5k 7.5k 76.4 37.1 57.5 65.8 71.3 87.6 76.7 86.9 69.91
ThinkLite-VL-72B 7.5k 79.7 43.8 64.3 68.3 72.0 88.2 77.3 87.7 72.67
$\Delta$ (Ours - Random selection) +3.3 +6.7 +6.8 +2.5 +0.7 +0.6 +0.6 +0.8 +3.06
$\Delta$ (Ours - Open 72B SoTA) +4.9 +8.6 +11.0 -2.0 +3.1 +0.8 +1.0 -0.2 +4.42
SoTA performance over both 7B and 72B models
@@
As shown in Table 2, ThinkLite-VL-7B and ThinkLite-VL-72B show a significant improvement in average performance across the eight benchmarks compared to the base model Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct, with the average performance increasing from 59.69 to 63.89 and 68.25 to 72.67, respectively. ThinkLite-VL-7B also outperforms reasoning models that primarily achieve performance enhancement through extensive knowledge distillation (such as LLaVA-CoT-11B, Mulberry-7B, Vision-R1-7B, and OpenVLThinker-7B) with the closest average performance to GPT-4o. Compared to MM-EUREKA-Qwen-7B, which does not involve SFT knowledge distillation but adopts a larger RL training dataset, our model consistently outperforms across all benchmarks, highlighting the importance of high-quality data filtering before training, and the effectiveness of the proposed MCTS-based filtering. For more discussion between offline and online data filtration, please refer to Section 4.3. Analyzing individual benchmarks, ThinkLite-VL-7B achieves best performance among all 7B-scale models on six out of eight benchmarks, with only marginal gaps behind InternVL2.5-7B on MMBench and MM-Vet. In addition, ThinkLite-VL-72B outperforms all existing open-source vision-language models across six benchmarks. Notably, ThinkLite-VL-7B attains SoTA accuracy of 75.1 on MathVista, exceeding both GPT-4o and o1. ThinkLite-VL-72B further advances the frontier, reaching 79.7 on MathVista and 64.3 on MathVerse, establishing new SoTA on both benchmarks.
Effectiveness of MCTS-based sample selection
Compared to training on an equal number of randomly selected samples from the full 70K dataset (ThinkLite-VL-7B-Random11k and ThinkLite-VL-72B-Random7.5k), ThinkLite-VL-7B and ThinkLite-VL-72B demonstrate a clear advantage across eight benchmarks, with average performance improvements of 5.4% at the 7B scale and 4.4% at the 72B scale. These results further show the importance of MCTS-based sample selection.
@@
4.2 Importance of MCTS-based Sample Selection
We conduct ablation studies to demonstrate the importance of MCTS-based sample selection. We compare five different training settings of ThinkLite-VL: (1) ThinkLite-VL-Unsolved: Trained using only the 5.6k samples that could not be solved by MCTS, representing the most difficult subset. (2) ThinkLite-VL-Iter5Only: Trained on the subset of data that VLM is able to solve via MCTS, but required more than 5 iterations. This set, combined with the unsolved samples, forms the full 11k training set used in ThinkLite-VL. (3) ThinkLite-VL-Random11k: Trained on a randomly sampled 11k subset from the full 70k dataset, matching the size of the ThinkLite-VL training set. (4) ThinkLite-VL-SelfConsistency: Trained on 23k samples selected based on a self-consistency difficulty measure. Specifically, for each prompt, we perform 50 rollouts using Qwen2.5-VL-7B-Instruct and compute answer accuracy using Qwen2.5-7B-Instruct. Samples with accuracy lower than 0.2 are selected for RFT. (5) ThinkLite-VL-Fullset: Trained on the complete 70k dataset without any filtering. We report the evaluation results of all five settings across the eight VLM benchmarks, as shown in Table 3.
We observe that ThinkLite-VL-7B, trained using 11k samples via MCTS-guided sample selection, achieves the highest average performance among all settings. It outperforms not only the random sampling baseline but also models trained on the full dataset and self-consistency-based filtering, despite using significantly fewer training samples. This highlights the effectiveness of our difficulty-aware data selection strategy. Further analysis reveals that models trained on subsets derived solely from unsolved samples or samples requiring more than five iterations also show decent performance, suggesting that hard and medium-difficulty samples contribute meaningfully to reasoning ability. However, neither subset alone is sufficient. The combination of both unsolved and medium-difficulty samples yields the strongest and most effective training signal. Additional analyses are in Appendix B.
4.3 Comparison with Online Data Selection
In this section, we compare our offline data-selection strategy with an online alternative and evaluate their impact on model performance. We adopt an online baseline based on self-consistency filtering: during training we keep only those samples whose rollout accuracy is greater than 0 but below 0.9, drawing additional samples until the training batch is full. Table 4 compares this online variant with our MCTS-based offline selector and a plain offline self-consistency baseline. Similar to the findings in other RL studies [90], the online filter offers negligible improvement except converges faster. The decisive factor is still the ability to identify examples that are truly challenging for the current model, a task at which our MCTS selector excels due to its explicit tree search.
Table 4: Comparison between ThinkLite-VL and model trained with offline and online self-consistency based sample selection. Our method demonstrates significant advantages.
Model Size Training type Selection method
MathVista
testmini
MathVision
mini
MathVerse
mini
MMMU
MMStar
MMBench
MM-Vet
AI2D
Avg.
7B Offline MCTS (Ours) 75.1 32.9 52.1 55.5 65.0 81.4 67.8 83.6 64.18
SelfConsistency 74.6 30.9 50.1 53.8 64.1 81.3 67.1 83.3 63.15
Online SelfConsistency 74.2 26.9 50.1 50.6 64.8 82.0 67.1 83.0 62.34
72B Offline MCTS (Ours) 79.7 43.8 64.3 68.3 72.0 88.2 77.3 87.7 72.67
SelfConsistency 77.3 39.1 62.0 66.3 71.6 87.7 77.0 87.1 71.01
Online SelfConsistency 76.9 38.5 58.2 66.0 71.7 87.5 77.1 87.4 70.12
4.4 Data Difficulty Analysis between 7B and 72B Models
We analyze the 11k and 7.5k sample sets selected by 7B and 72B models, to examine how models of different capacity agree on the sample difficulty. We find that there is an overlap of 5.4k samples, where 3.6k of them are instances that neither model is able to solve within 50 MCTS iterations. The real divergence lies in the mid-difficulty stratum. We observe that for this subset, the two models often behave asymmetrically: problems easily solved by the 7B model may require many more iterations for the 72B model, and vice versa, exposing distinct reasoning heuristics across models.
We validate this model-specific preference through cross-sample training: the 11k samples selected by the 7B model are used to RFT the 72B model, and vice versa. Table 5 shows that the gains in both settings were markedly smaller than when each model trains on its own curated set. These results suggest that a sample set tailored to one model transfers poorly to another, even in a strong-to-weak setting. Instead, it is more effective to scale extra compute to find appropriately difficult samples that best fit the model itself, as the approach proposed in ThinkLite-VL.
Table 5: Comparison between the 7B and 72B models which trained on each other’s selected samples, the resulting performance improvements drops significantly.
Models Data size
MathVista
testmini
MathVision
mini
MathVerse
mini
MMMU
MMStar
MMBench
MM-Vet
AI2D
Avg.
ThinkLite-VL-7B 7.5k-72B 70.2 26.3 49.2 51.6 61.7 81.1 66.9 82.9 61.24
11k-7B 75.1 32.9 52.1 55.5 65.0 81.4 67.8 83.6 64.18
ThinkLite-VL-72B 11k-7B 76.4 38.5 58.4 67.2 70.2 87.3 76.6 87.4 70.24
7.5k-72B 79.7 43.8 64.3 68.3 72.0 88.2 77.3 87.7 72.67
5 Conclusion
We have introduced an effective self-improvement approach to enhance the reasoning capabilities of VLMs, eliminating the need for external supervision or knowledge distillation. Our key insight highlights the critical importance of selecting appropriately challenging examples for RFT. We find that when training data quality is sufficiently high, even a small dataset can substantially enhance visual reasoning performance without knowledge distillation. Building on this insight, we propose a novel data selection technique, MCTS-based sample selection, which identifies and retains challenging samples by quantifying the number of MCTS reasoning iterations. Starting from 70k initial samples, we obtain a high-quality subset comprising 11k and 7.5k challenging samples for 7B-level and 72B-level models, respectively. These curated datasets are then used to fine-tune the Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct via RFT, resulting in the reasoning VLMs named ThinkLite-VL-7B and ThinkLite-VL-72B. Our models demonstrate significant improvements across multiple visual reasoning benchmarks, and notably achieves a new SoTA accuracy of 79.7 on MathVista and 64.3 on MathVerse. We hope that our findings on the difficulty-based selection of RFT training data can provide insights for training more effective reasoning VLMs.
Acknowledgment
Appendix A Prompts used in experiments
A.1 Prompt for MCTS
The prompt used for MCTS is shown in Table 6.
[…]
A.2 Prompt for Critic Model
The prompt used for critic model during MCTS is shown in Table 7.
Table 7: Critic prompt for MCTS simulation results evaluation.
A.3 Prompt for RFT
The prompt used for RFT is shown in Table 8.
[…]
Appendix B More experiments
B.1 Reward curves of VLM with different training data
We compare the reward curves during RFT of ThinkLite-VL-Random11k, ThinkLite-VL-Fullset, ThinkLite-VL-Iter5Only, and ThinkLite-VL, as shown in Figure 5. Although ThinkLite-VL-Random11k and ThinkLite-VL-Fullset achieve higher rewards during training, their actual benchmark performances are inferior to ThinkLite-VL. This observation suggests that incorporating a large number of easy samples into training rapidly improves rewards but fails to enhance the model’s reasoning ability. Moreover, ThinkLite-VL exhibits notably lower rewards compared to ThinkLite-VL-Iter5Only, indicating that the unsolved data identified by our MCTS-based sample selection strategy indeed pose significant challenges to the VLM. By progressively learning to solve these challenging problems during training—even if not all are solved completely—the reasoning capabilities of VLMs can be substantially improved.
@@ B.1 Reward curves of VLM with different training data
Figure 5: Comparison of reward curves of 7B models trained with different data during RFT. Iter5+Unsolved 11k dataset presents the most challenging learning setting for VLM, highlighting the difficulty of the samples selected by MCTS-based sample selection.
B.2 Ablation Study of Data Difficulty
@@
Table 9: ThinkLite-VL-7B performance under different training data difficulty settings. Iter5+Unsolved achieves the best performance.
@@
Appendix C Case Studies
In this section, we present samples of varying difficulty levels selected by the MCTS-based sample selection method across different datasets, as shown in Tables 15 through 14. The difficulty levels are determined based on the number of reasoning iterations required by the VLM to arrive at the correct answer during the MCTS process, providing reference examples for understanding how the method distinguishes between easy and challenging samples.